If you were supposed to use a disinfectant solution in a1 to 40 ratio of concentrate to water how much concentrate should you add to 5 liters of water
Answer by stanbon(75887)

If you were supposed to use a disinfectant solution in at 1 to 40 ratio of concentrate to water how much concentrate should you add to 5 liters of water
---
concentract/water = 1/40
----
c/5 = 1/40
---
c = 5/40
---
c = 1/8 liter
-----------
Cheers,
Stan H.
-------------